# Bridged Semantic Alignment for Zero-shot 3D Medical Image Diagnosis11footnotemark: 1

### v1

@@ 5.1 Ablation Study
5.2 Hyper-parameter of CMKB
@@
The hyper-parameter $K$ in CMKB, which defines the number of basis vectors, plays a pivotal role in determining its spatial representational capacity. To evaluate the influence of $K$, we conduct experiments with values $K$ = {1024, 2048, 3072, 4096}. As illustrated in Fig. 6, increasing $K$ enhances the spatial coverage of CMKB, enabling the model to capture a broader range of features and improve image-text alignment quality. However, this improvement diminishes and eventually saturates as $K$ becomes excessively large. In such scenarios, an overabundance of basis vectors can hinder cross-modal information interaction. For instance, when $K$ is excessively large, the model may reconstruct features independently within each modality, effectively bypassing the role of CMKI and nullifying cross-modal interaction. Based on these observations, we set $K=2048$ for all subsequent experiments, as it strikes an optimal balance between alignment performance and computational efficiency.
@@
Fig. 8: Generic descriptions of chest CT generated by a LLM for visualizing text features.
@@ 5.3 Visualization of Features
Fig. 1 and Fig. 7 illustrate the UMAP visualization of image, text, and CMKB features. To assess the effectiveness of vision-language alignment, we ensure that all text descriptions are aligned with all CT volumes. As shown in Fig. 8, we leverage GPT-4 to generate multiple unbiased, generic descriptions of chest CT volumes as text samples for feature extraction. During the UMAP computation, cosine similarity is employed to measure the distances between different features, representing the degree of alignment among them. To quantify the impact of introducing CMKI on vision-language alignment, we define the Feature Alignment Score (FAS) as:
[…]
where the silhouette score, originally used to evaluate cluster separation [35, 37, 38], measures the cohesion of image and text feature sets. By normalizing the silhouette score to the range of [0, 1], the FAS quantifies the alignment: a score closer to 1 indicates better alignment, while a score closer to 0 indicates poorer alignment.
Comparing Fig. 1 (b) and Fig. 7 (a), we observe that our proposed BrgSA framework significantly reduces the gap between image and text feature spaces, resulting in higher FAS values (0.37 vs. 0.50). This demonstrates that CMKI effectively enhances vision-language alignment. Further, Fig. 7 (b) and Fig. 7 (c) reveal that CMKB features align more closely with image features than with text features. This discrepancy can be attributed to the high-level abstraction and relatively uniform nature of text features, whereas image features are richer and more diverse. The greater variability in image feature representation makes CMKB features gravitate toward image features. Finally, as shown in Fig. 1 (c), CMKB features act as a bridge between image and text features, connecting the two feature spaces and facilitating their interaction. This bridging role effectively improves the overall vision-language alignment, further validating the importance of CMKB in this process.
Fig. 9: Comparison of mean differences between paired and unpaired image-text samples on CT-RATE validation dataset.
@@ 5.4 Mean Difference for Pair-unpair Discrimination
5.5 Limitations and future work
Despite the promising results achieved by our proposed method, several limitations remain to be addressed. First, while our method incorporates semantic summarization and cross-modal knowledge interaction for global feature alignment, it may still overlook fine-grained local details in 3D images, which are critical for certain complex abnormalities. Second, while the zero-shot setting demonstrates strong performance, further exploration of few-shot scenarios may provide more comprehensive insights into the method’s adaptability and robustness.
In future work, we aim to address these limitations by enhancing the integration of fine-grained local feature alignment into our framework to better capture subtle details in 3D images. Additionally, extending the current approach to include few-shot learning settings and exploring its application in broader clinical scenarios will be the focus of our ongoing research.
@@ 6 Conclusion
Data availability
The code and model will be made publicly available online. The dataset used is publicly accessible online. Additionally, the custom annotations will also be made available online.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used GPT-4 from OpenAI in order to improve readability and language. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
Acknowledgments
This work was supported by Natural Science Foundation of China under Grant 62271465, Suzhou Basic Research Program under Grant SYG202338, and Open Fund Project of Guangdong Academy of Medical Sciences, China (No. YKY-KF202206).
Supplementary material
Supplementary material is prepared and provided as a separate file along with the manuscript.
Supplementary Material
As shown in Table S1S4, we provide a comprehensive benchmark of different methods on the CT-RATE-LT dataset. Specifically, we present the detailed performance metrics for each abnormality, including Area under the ROC Curve (AUC), Accuracy, F1 score and Precision. These results enable a direct comparison of the effectiveness of various methods in detecting specific abnormalities, offering valuable insights into their strengths and limitations.
Table S1: AUC comparison of different methods for various abnormalities on CT-RATE-LT. The best performance is highlighted as bold.
Method PN ATH SPD PTX BT OST EMP LC LM PC PMD PF TB PLC RF Mean
VocabFine 66.73 80.94 60.45 61.80 36.40 99.36 96.48 54.08 67.76 97.14 74.99 6.32 67.34 85.25 82.69 69.18
CT-CLIP 63.99 78.46 65.32 74.15 63.78 76.70 93.25 64.19 84.41 77.66 90.29 94.11 91.16 81.93 54.51 76.93
Merlin 59.85 84.04 59.40 84.54 68.42 59.28 95.34 77.54 67.06 79.34 31.89 84.71 90.32 92.76 95.56 75.34
BrgSA 79.51 87.05 84.16 98.50 65.75 97.54 97.28 83.26 84.15 79.01 74.93 73.71 95.46 84.43 99.51 85.62
Note: PN: Pneumonia, ATH: Atherosclerosis, SPD: Spondylosis, PTX: Pneumothorax, BT: Bronchial thickening, OST: Osteophyte, EMP: Empyema, LC: Lung cyst, LM: Lung mass, PC: Pericardial calcification, PMD: Pneumomediastinum, PF: Pulmonary fibrosis, TB: Tuberculosis, PLC: Pleural calcification, RF: Rib fracture.
Table S2: Accuracy comparison of different methods for various abnormalities on CT-RATE-LT.
Method PN ATH SPD PTX BT OST EMP LC LM PC PMD PF TB PLC RF Mean
VocabFine 85.46 98.65 87.33 46.36 44.72 99.70 95.82 47.58 62.32 96.38 71.67 5.82 58.54 84.76 82.43 71.17
CT-CLIP 86.97 93.85 67.62 81.34 45.87 80.85 91.25 59.20 79.99 76.21 86.57 92.79 93.78 81.80 53.97 78.14
Merlin 81.90 99.34 49.29 99.64 96.05 44.52 93.45 91.94 89.93 88.32 12.90 82.23 95.33 92.66 95.06 80.84
BrgSA 89.60 98.88 97.43 98.91 83.84 99.47 96.78 80.29 83.19 76.54 74.10 70.25 94.37 84.37 99.47 88.50
Table S3: F1 score comparison of different methods for various abnormalities on CT-RATE-LT.
Method PN ATH SPD PTX BT OST EMP LC LM PC PMD PF TB PLC RF Mean
VocabFine 89.08 98.87 92.99 63.13 61.65 99.77 97.80 64.41 76.72 98.09 83.43 10.89 73.78 91.72 90.33 79.51
CT-CLIP 89.90 96.36 80.43 89.53 62.71 89.32 95.36 74.30 88.82 86.43 92.74 96.20 96.73 89.96 70.06 86.59
Merlin 86.97 99.22 65.76 99.65 97.83 61.50 96.55 95.74 94.63 93.73 22.75 90.18 97.54 96.16 97.44 86.38
BrgSA 91.78 99.00 98.46 99.30 91.06 99.66 98.30 89.00 90.76 86.64 85.06 82.46 97.04 91.49 99.71 93.31
Table S4: Precision comparison of different methods for various abnormalities on CT-RATE-LT.
Method PN ATH SPD PTX BT OST EMP LC LM PC PMD PF TB PLC RF Mean
VocabFine 7.95 6.67 0.52 0.37 0.18 12.50 1.55 0.13 0.17 1.79 0.23 0.07 0.16 0.22 0.19 2.18
CT-CLIP 7.53 2.22 0.61 0.53 0.30 0.34 0.75 0.16 0.33 0.28 0.49 0.90 0.53 0.18 0.07 1.01
Merlin 6.92 14.29 0.45 14.29 0.85 0.18 1.00 0.41 0.33 0.28 0.08 0.37 0.70 0.45 0.66 2.75
BrgSA 16.56 8.70 5.13 11.43 0.41 11.76 2.00 0.33 0.39 0.28 0.25 0.22 1.16 0.21 5.88 4.31

### v2

@@ 5.4 Ablation Study
Table 6: Ablation study on the Semantic Summarization ($S$) and the CMKI module for zero-shot abnormality diagnosis on the RAD-ChestCT dataset.
$S$ CMKI AUC ACC F1 Prec mAP Rec@1 Prec@3
65.0 61.1 65.4 35.7 39.3 12.9 38.9
67.9 64.0 67.9 37.2 40.9 13.3 42.2
68.0 64.0 67.9 38.0 41.7 12.3 40.2
70.0 65.5 69.3 39.1 43.0 14.3 44.0
Table 7: Ablation study on the impact of the Semantic Summarization ($S$) and the CMKI module for the report-to-volume retrieval task on the CT-RATE validation dataset.
$S$ CMKI Recall@5 Recall@10 Recall@50 Recall@100
2.2 3.7 13.1 22.2
2.2 4.1 18.4 29.1
3.4 5.8 21.8 34.6
5.8 10.1 28.6 42.0
5.5 Robustness of Semantic Summarization
To evaluate the reliability of semantic summarization in our framework, we conduct two experiments to assess the reliability of semantic summarization in our framework.
(a) Summarizer comparison. We compare summarizers of different scales as well as domains, including general-domain models (e.g., Qwen-2.5, Llama-3, GPT-4, DeepSeek) and medical-domain models (e.g., Lingshu, Medgemma).
(b) Report perturbations. We introduce three types of noise before summarization: (1) style conversion with shortened and lengthened expressions, (2) grammatical errors in five sentences per report, and (3) random removal of 30% or 40% of sentences. The perturbed reports are then summarized using semantic summarization module based on GPT-4 for downstream alignment and zero-shot classification.
Results in Table 8 reveal three consistent trends: (1) models of different scales show comparable summarization quality, and larger models do not consistently outperform smaller ones; (2) medical-domain models provide performance on par with general-domain counterparts, indicating limited domain-specific advantage; and (3) the pipeline remains robust under report perturbations, maintaining competitive accuracy even with grammatical errors or missing sentences. These findings demonstrate that our approach is robust and generalizable, even in noisy or imperfect clinical reporting scenarios.
Table 8: Zero-shot diagnostic performance under different summarizers (a) and under report perturbations (b).
(a) Summarizer Comparison on Original Reports
Summarizer AUC ACC F1 Prec mAP Rec@1 Prec@3
Smaller-scale Language Models
Qwen-2.5-7B [46] 69.1 64.6 68.5 38.1 41.8 15.3 45.3
Lingshu-7B [47] 69.2 64.9 68.7 38.3 42.4 9.9 43.0
Llama-3-8B [48] 68.6 65.1 68.9 38.4 42.0 12.3 44.3
Medgemma-4b [49] 68.6 64.7 68.5 37.8 41.7 10.8 43.7
Larger-scale Language Models
Medgemma-27b [49] 69.2 64.8 68.6 38.1 42.5 14.8 43.2
Deepseek-V3 [50] 68.6 65.9 69.5 38.5 42.6 13.5 42.9
GPT-4 Turbo 70.0 65.5 69.3 39.1 43.0 14.3 44.0
(b) Robustness Under Report Perturbations
Perturbation AUC ACC F1 Prec mAP Rec@1 Prec@3
Baseline 70.0 65.5 69.3 39.1 43.0 14.3 44.0
Shortened expressions 69.9 65.8 69.5 38.9 43.3 14.0 43.7
Lengthened expressions 68.3 64.6 68.5 37.8 41.9 15.4 45.8
Grammatical errors 67.6 63.6 67.5 37.5 41.1 13.5 43.0
Missing sentences (30%) 67.4 64.5 68.3 37.6 41.1 10.3 40.4
Missing sentences (40%) 66.9 64.4 68.2 37.7 41.4 9.6 39.2
Figure 6: Ablation results on the RAD-ChestCT dataset, showing the effect of varying $K$ in CMKB as well as the impact of the hyperparameters $\alpha$ and $\gamma$ in the total alignment loss.
Table 9: Ablation study of CMKB initialization strategies, including random distributions and pretrained embeddings, evaluated on RAD-ChestCT for zero-shot diagnosis.
CMKB Initialization AUC ACC F1 Prec mAP Rec@1 Prec@3
Random initialization
Gaussian Random 68.7 64.6 68.4 38.3 41.6 13.8 43.4
Xavier Random 68.8 64.7 68.6 38.2 42.0 14.8 43.0
Uniform Random 68.5 64.2 68.1 37.6 41.6 13.7 43.1
Orthogonal Random 70.0 65.5 69.3 39.1 43.0 14.3 44.0
Embedding-based initialization
Image Embedding 68.4 64.4 68.3 38.3 42.4 10.6 43.7
Text Embedding 63.9 61.7 65.8 35.4 37.5 11.7 40.4
Cross Embedding 65.4 61.9 66.0 36.0 39.1 14.1 42.0
UMLS [51] 68.7 64.5 68.4 37.9 41.8 14.1 42.2
RadLex [52] 69.4 65.5 69.2 38.7 43.3 14.0 43.1
5.6 Effect of Initialization Strategies in CMKB
We evaluate the effect of different initialization schemes for CMKB, comparing random distributions and embedding-based priors. Random strategies include Gaussian, Xavier, Uniform, and Orthogonal initialization. For embedding-based strategies, we use features extracted from the vision and text encoders pretrained with M3AE. Image initialization applies $k$-means clustering on vision encoder features from the training set, while Text uses text encoder features. Cross combines both modalities before clustering to form cross-modal centroids. In addition, UMLS [51] and RadLex [52] are treated as textual corpora, encoded with the text encoder and clustered into pretrained embeddings. As shown in Table 9, orthogonal random initialization yields the best results among random schemes, whereas RadLex provides the best embedding-based initialization, demonstrating the benefit of domain-specific ontologies for CMKB. The relatively poor performance of Text initialization may be due to the repetitive and template-driven nature of training reports, which could lead to clustered embeddings lacking diversity.
5.7 Hyperparameter Sensitivity Analysis
@@
Effect of $K$. The hyper-parameter $K$ in CMKB, which defines the number of basis vectors, plays a pivotal role in determining its spatial representational capacity. To evaluate the influence of $K$, we conduct experiments with values $K$ = {1024, 2048, 3072, 4096}. As illustrated in Fig. 6, increasing $K$ enhances the spatial coverage of CMKB, enabling the model to capture a broader range of features and improve image-text alignment quality. However, this improvement diminishes and eventually saturates as $K$ becomes excessively large. In such scenarios, an overabundance of basis vectors can hinder cross-modal information interaction. For instance, when $K$ is excessively large, the model may reconstruct features independently within each modality, effectively bypassing the role of CMKI and nullifying cross-modal interaction. Based on these observations, we set $K=2048$ for all subsequent experiments, as it strikes an optimal balance between alignment performance and computational efficiency.
Effect of loss weights. We further study the loss hyper-parameters $\alpha$ and $\gamma$, with $\beta$ fixed to 1. As illustrated in Fig. 6, moderate values of $\alpha$ (around 0.5) enhance performance by strengthening reconstruction constraints, while larger values lead to performance degradation. Setting $\alpha=0$ markedly reduces both AUC and F1, confirming that the reconstruction loss $\ell_{\text{MSE}}$ is essential. Similarly, $\gamma$ controls the contribution of the InfoNCE on reconstructed features $\ell_{\text{INFO-R}}$. Small to moderate values improve alignment, with the best results obtained near $\gamma=1$. Removing this term ($\gamma=0$) results in weaker alignment and lower performance, demonstrating its necessity. For completeness, we also examined the case of removing the primary contrastive alignment loss by setting $\beta=0$ (i.e., ablating $\ell_{\text{INFO}}$). This caused a drastic performance drop (AUC 51.8, F1 54.3), far below all other variants, clearly indicating that $\ell_{\text{INFO}}$ is indispensable for maintaining alignment quality.
Effect of batch size. We also examine the influence of batch size on InfoNCE optimization, as shown in Fig. 7. Larger batch sizes consistently yield better performance: for example, the AUC improves from 70.0 at batch size 8 to 72.6 at batch size 64, while the F1 score rises from 69.3 to 71.0. This aligns with the intuition from CLIP training that larger batches provide more diverse negative samples within each iteration, strengthening contrastive alignment between modalities. Although hardware resources limit the maximum batch size in our experiments, future work could explore scaling to larger batches for further gains.
Figure 7: Results of varying batch size for BrgSA on RAD-ChestCT dataset.
Figure 8: UMAP visualization of features. Cosine similarity is used to evaluate the alignment between image and text features. The text features are generated using generic descriptive texts to ensure that all images can be matched to all texts. (a)–(c) represent features trained by BrgSA.
Table 10: Ablation on vision and text encoders for CLIP-based zero-shot diagnosis. We control one encoder while varying the other to ensure fair single-variable comparison.
(a) Vision Encoder Exploration (Fixed Text: PubMedBERT)
Vision Encoder AUC ACC F1 Prec mAP Rec@1 Prec@3
ResNet18 [53] 70.8 66.6 70.2 39.7 45.2 14.6 46.2
ResNet34 [53] 70.4 66.3 69.9 39.7 43.6 13.0 45.6
ResNet50 [53] 70.4 66.0 69.8 39.4 43.6 14.7 46.8
SwinVit-B [54] 72.2 67.7 71.1 40.5 47.3 13.8 47.1
ViT-B [39] 72.6 67.7 71.1 40.9 47.6 14.9 46.8
(b) Text Encoder Exploration (Fixed Vision: ViT-B/16)
Text Encoder AUC ACC F1 Prec mAP Rec@1 Prec@3
BioClinicalBERT [55] 72.7 68.0 71.4 41.6 47.5 13.8 46.0
CXRBERT [40] 74.2 68.6 72.0 42.2 49.4 14.9 48.7
BioGPT [56] 64.9 62.8 66.9 36.2 38.7 13.0 42.2
PubMedBERT [57] 72.6 67.7 71.1 40.9 47.6 14.3 46.8
5.8 Exploration of Vision and Text Encoders
Table 10 shows the impact of different vision and text backbones. For vision encoders, ViT-B performs best, benefiting from global self-attention that captures long-range dependencies in CT volumes. For text encoders, while we initially used PubMedBERT [57], CXRBERT [40] achieves the best results due to its pretraining on radiology reports. BioClinicalBERT is competitive but less radiology-specific, and BioGPT underperforms as a generative model. Overall, ViT-B and CXRBERT prove most effective, and we adopt them as the final encoders in our framework.
5.9 Inference Efficiency Benchmark
We benchmark diagnostic performance and efficiency on RAD-ChestCT (Table 11). Our method achieves the highest AUC (74.2) with comparable latency (43.1 ms/study) and memory usage (1.7 GB), clearly outperforming prior approaches in accuracy. While fVLM [15] shows slightly lower latency and GPU memory, its pipeline requires an additional automatic segmentation step whose computational cost is not included in the table, making its reported efficiency less favorable in practice. Furthermore, by directly running inference with reduced input resolutions, our model maintains competitive performance (AUC 67.7–72.2) while reducing latency to 21.7 ms and memory to 1.1 GB, highlighting its strong adaptability for real-world deployment.
Table 11: Evaluation on RAD-ChestCT: AUC with inference efficiency (time per study and peak GPU memory).
Method Height Width Slices AUC Time (ms/study) $\downarrow$ GPU Mem (GB) $\downarrow$
[…]
@@ 5.9 Inference Efficiency Benchmark
Figure 9: Generic descriptions of chest CT generated by a LLM for visualizing text features.
@@ 5.10 Visualization of Features
Fig. 1 and Fig. 8 illustrate the UMAP visualization of image, text, and CMKB features. To assess the effectiveness of vision-language alignment, we ensure that all text descriptions are aligned with all CT volumes. As shown in Fig. 9, we leverage GPT-4 to generate unbiased, generic descriptions of chest CT volumes as text samples for feature extraction. During the UMAP computation, cosine similarity is employed to measure the distances between different features, representing the degree of alignment among them. To quantify this alignment, we first introduce the silhouette score [58], originally used to evaluate cluster separation. For each sample $i$, let $a(i)$ denote the mean intra-cluster distance and $b(i)$ the mean nearest-cluster distance. The silhouette score is then defined as
$\color[rgb]{0,0,0}s(i)=\frac{b(i)-a(i)}{\max\{a(i),\,b(i)\}},\quad S=\frac{1}{N}\sum_{i=1}^{N}s(i).$} (12)
In our setting, the three feature sets (image, text, and CMKB) are predefined rather than obtained by unsupervised clustering. Hence, each sample has a fixed cluster assignment and no mis-clustering can occur. Consequently, the intra-cluster distance $a(i)$ is never larger than the inter-cluster distance $b(i)$, ensuring that $s(i)$ remains non-negative and $S$ lies in $[0,1]$. To adapt this metric for evaluating vision-language alignment, we define the Feature Alignment Score (FAS) as
$\color[rgb]{0,0,0}\text{FAS}=1-S,$} (13)
where a score closer to 1 indicates better alignment, while a score closer to 0 indicates poorer alignment.
Comparing Fig. 1 (b) and Fig. 8 (a), we observe that BrgSA significantly reduces the gap between image and text feature spaces, resulting in higher FAS values (0.37 vs. 0.50). This demonstrates that CMKI effectively enhances vision-language alignment. Further, Fig. 8 (b) and Fig. 8 (c) reveal that CMKB features align more closely with image features than with text features. This discrepancy can be attributed to the high-level abstraction and relatively uniform nature of text features, whereas image features are richer and more diverse. The greater variability in image feature representation makes CMKB features gravitate toward image features. Finally, as shown in Fig. 1 (c), CMKB features act as a bridge between image and text features, connecting the two feature spaces and facilitating their interaction. This bridging role effectively improves the overall vision-language alignment, further validating the importance of CMKB in this process.
Figure 10: Comparison of mean differences between paired and unpaired image-text samples on CT-RATE validation dataset.
@@ 5.11 Mean Difference for Pair-unpair Discrimination
Figure 11: Pairwise confusion analysis of Pneumothorax (PNX) and Atelectasis (ATX): $\times2$ XOR-subset (left) and $\times4$ joint states (right). Cells show row-normalized percentages with counts in parentheses.
6 Discussion
In this work, we present BrgSA, a bridged semantic alignment framework that closes the gap between 3D medical images and diagnostic texts for zero-shot disease recognition. BrgSA combines InfoNCE contrastive learning with CMKI-based masked cross-modal reconstruction to yield robust, interpretable alignment. We then discuss its clinical and technical impact, innovations, and limitations.
6.0.1 Clinical Impact
Zero-shot learning offers a promising solution for diagnosing unseen diseases, particularly rare conditions that have very limited annotated samples and are difficult to handle using conventional supervised methods. Our proposed method, BrgSA, enhances vision-language alignment in 3D medical data and significantly improves zero-shot diagnostic performance. This improvement is of great clinical relevance for rare disease diagnosis, where prior examples are scarce. To support clinical validation, we construct CT-RATE-LT, a benchmark dataset featuring underrepresented abnormalities. Our model achieves strong results on both CT-RATE-LT and the RadChest-LT benchmark, demonstrating its robustness across different datasets. Moreover, unlike many prior methods focusing on 2D images, our approach is designed for 3D CT, a clinically critical imaging modality, further underscoring its potential for real-world medical applications. We also provide confusion analyses of clinically related findings, such as pneumothorax versus atelectasis (Fig. 11). The results indicate that while the model can distinguish the two abnormalities in many cases, confusion remains—particularly with pneumothorax being misclassified as atelectasis and atelectasis often under-detected. These findings highlight both the potential and the limitations for clinical application.
6.0.2 Technical Impact
We propose a simple yet effective framework to enhance vision-language alignment in 3D medical data. At the core of our method propose the CMKB, which acts as a semantic bridge between image and text features. By enabling implicit interaction across modalities, CMKB reinforces shared semantics and alleviates modality-specific biases. This design leads to more accurate and robust feature alignment, which is critical for improving zero-shot diagnostic performance in 3D medical imaging tasks.
6.0.3 Innovation
This work introduces two key innovations to advance zero-shot 3D medical image diagnosis. First, we propose the CMKI module, which implicitly aligns 3D visual and textual features by bridging the modality gap. This component complements contrastive learning and enables more stable and semantically coherent alignment across modalities. Second, we construct CT-RATE-LT, a benchmark test set comprising underrepresented abnormalities, to evaluate zero-shot performance on rare and low-frequency conditions.
6.0.4 Limitations and Future Work
While our method demonstrates strong performance in zero-shot 3D medical diagnosis, several limitations remain. First, although the framework effectively integrates semantic summarization and cross-modal knowledge interaction for global alignment, it may still underrepresent fine-grained local features that are essential for detecting subtle or complex abnormalities in 3D volumes. Second, by contrasting success and failure cases, we find that the model occasionally misaligned its attention, overlooking true pathological regions and resulting in missed detections in low-confidence predictions (Fig. 12). Third, our current focus on zero-shot generalization leaves open the question of how well the method performs under limited supervision. Fourth, since the current evaluation is limited to thoracic CT datasets, the method’s generalizability to other anatomical domains remains to be validated. Fifth, as our framework partially relies on LLM-based semantic summarization, potential biases inherent in pretrained LLMs may influence the generated representations. Sixth, the framework’s relatively high computational cost could pose deployment challenges in resource-constrained clinical environments. To address these limitations, future work will enhance local feature alignment and attention calibration to improve sensitivity to subtle abnormalities. We also plan to investigate adaptive learning strategies for broader generalization, while exploring lightweight architectures and bias-mitigation techniques to enable fair and efficient deployment in real-world clinical settings.
Figure 12: Representative cases with attention heatmaps (top) and ground-truth masks (bottom). (a),(b) show successful predictions with high similarity scores for positive findings, while (c),(d) illustrate missed cases with low similarity despite positive ground truth.